15. In constructing a study of the most popular car colors in a city the researcher stood on a corner and counted the autos of each color that pass by. This might not be the best way to do this because a. the answer would depend on the corner b. the answer would depend on the time of day c. the answe would depend on the frequency car owners of each color drive their cars d. all of the above 16. A double blind study is especially useful when a.) the issues are numerous and unclear b.) when the experimenter has a personal interest in the result c.) when the subjects in the study have eye problems d.jnone of the above 17. In a small weekly lottery there are 1,000 tickets sold and three of them are always winners. What is the probability that George has winning ticket in three consecutive lottery? a. 1/1000. b. 3/1000. c. (1/1000). d. (3/1000) 18. A certain US Senator is a devoted family man and lives modestly. Given these facts, which is more likely de took a bribe from a lobbyist B. He took a bribe from a lobbyist and used it to pay for his daughters medical needs. B) c. equal d. cant tell
Answer #1

Answer:

15. All options are correct(All the above is the answer)

16. All are rong (None of the above is correct answer)

17. =(3 / 1000) * (3/1000) * (3/1000) = (3/1000)^3